OpenStudy (anonymous):

You deposit $100 into an account that pays 6% annual interest. How long will it take for the balance to reach 1000 if the interest is compounded with the given frequency? A. Quarterly B. Continuously

OpenStudy (wolf1728):

6% annually = (1+.06/4)^4 = 1.0613635506 6.13635506 % compounded quarterly When it is continuously compounded, the formula is annual = e^r -1 annual = (2.718281828 ^ .06) -1 annual = 1.0618365465 -1 annual rate = .0618365465 or 6.18365465% To see those formulas: http://1728.org/rate.htm Now to do the other calcuations.

OpenStudy (wolf1728):

We need to know how long it takes $100 to become $1,000 at 6.13635506 % The formula is: Years = [log (total) - log(principal)] / log (1 + rate) Years = (log 1,000 - log 100) / log (1.0613635506) Years = (3 -2) / 0.025864169 Years = 1 / 0.025864169 Years = 38.6635271566 Now for the continuously compounded interest: $100 to become $1,000 at 6.18365465 % The formula is: Years = [log (total) - log(principal)] / log (1 + rate) (except for the rate, this is the same as above) Years = (3-2) / log (1.0618365465) Years = 1 / 0.0260576689 Years = 38.3764182375 To see those formulas: http://1728.org/compint2.htm http://1728.org/rate.htm